### v1

# The Dance of the Sheared Eigenfunctions

Abstract
In this work, we delve into the theory of sheared potentials in non-relativistic quantum mechanics. After defining what we mean by a family of sheared potentials, we consider these families in two particular but emblematic cases, the harmonic oscillator and the symmetric potential well proportional to $|x|$. In both cases, besides determining the spectra, we analyse the impact of the shearing process on the respective eigenfunctions. The latter analysis is typically left aside in the literature, but here we show that the sheared eigenfunctions yield insights that allow for a deeper understanding of the main features exhibited by the spectra. Finally, we make a few comments about the connection between the change in the spectra of a given sheared family and the necessary work that must be made by an external agent to implement such a change.
  • April 2025
1 Introduction: sheared potentials in quantum theory
Potential theories are one of the most important ingredients of a physicist toolbox. Most equations in physics are, in fact, schemes of differential equations with distinct potential functions, as for instance, the Laplace and the Schrödinger equations. Naturally, solving these equations in the widest possible generality is a well established goal in many branches of physics. With this purpose, we consider an interesting and yet underdeveloped method of creating a parametrized family of potentials which are particularly appealing to be explored in the Schrödinger equation, namely, sheared potentials.
In order to define precisely what we mean by sheared potentials, let us consider one dimensional periodic motions of a particle along a given ${\cal O}x$ axis, which is submitted to a conservative force $F(x)=-\mathcal{U}^{\prime}(x)$ where $\mathcal{U}(x)$ is the associated potential. For periodic motions this must be a well of some sort, and we assume it to have just one minimum, which we take to be at $x=0$. Let $\mathcal{E}$ denote the mechanical energy measured with respect to the minimum potential energy and also $a(\mathcal{E}),b(\mathcal{E})$ where $a(\mathcal{E})<0<b(\mathcal{E})$ be the turning points of the motion, determined by the roots of the equation $\mathcal{E}=\mathcal{U}(x)$.
We define a family of sheared potential wells by the set of potentials $\{\mathcal{U}_{\xi}\}$, with $\xi$ being a continuous parameter which deforms continuously the original function $\mathcal{U}$, constrained by the following condition
$b_{\xi^{\prime}}(\mathcal{E})-a_{\xi^{\prime}}(\mathcal{E})=b_{\xi}(\mathcal{E})-a_{\xi}(\mathcal{E})\,,$ (1)
where $a_{\xi}(\mathcal{E}),b_{\xi}(\mathcal{E})$ are the turning points of $\mathcal{U}_{\xi}(x)$ for the mechanical energy $\mathcal{E}$ (analogously for $a_{\xi^{\prime}}(\mathcal{E}),b_{\xi^{\prime}}(\mathcal{E})$), for any allowed value of ${\cal E}$. Equation (1) defines a family of sheared potentials, and without loss of generality we assume that in the deformation process the minimum value of $\mathcal{U}_{\xi}$ is zero and happens at $x=0$.
In Fig. 1 we plot two potential wells which belong to the same family of sheared potentials. The shearing condition given by Eq. (1) can be understood graphically in this figure with the aid of the double horizontal arrows.
$a_{\xi}$$b_{\xi}$$a_{\xi^{\prime}}$$b_{\xi^{\prime}}$$\mathcal{E}$$\mathcal{U}_{\xi}(x)$$\mathcal{U}_{\xi^{\prime}}(x)$
Figure 1: Two potential wells belonging to the same family of sheared potentials. Note that the shearing condition, $b_{\xi^{\prime}}-a_{\xi^{\prime}}=b_{\xi}-a_{\xi}$, is fullfilled.
Classically, families of sheared potential wells are very rich because it can be shown that the periodic motions associated to potential wells belonging to the same family have the same period (as a function of the mechanical energy) [1]:
$\tau_{\xi}(\mathcal{E})=\sqrt{2m}\int_{a_{\xi}}^{b_{\xi}}\dfrac{dx}{\sqrt{\mathcal{E}-\mathcal{U}_{\xi}(x)}}=\sqrt{2m}\int_{a_{\xi^{\prime}}}^{b_{\xi^{\prime}}}\dfrac{dx}{\sqrt{\mathcal{E}-\mathcal{U}_{\xi^{\prime}}(x)}}=\tau_{\xi^{\prime}}(\mathcal{E})\,,$ (2)
for all values of $\xi$ and $\xi^{\prime}$ of the same family. The study of sheared potentials is well established on the classical regime [2, 4, 3, 5], but it is not so widespread in the quantum realm. The investigation of sheared potentials is of interest for the Einstein-Brillouin-Keller quantization scheme and for isocanonical ensembles in statistical mechanics [6], as also for the study of Cherkas systems [7], superintegrability of Hamiltonians [8] and the general modern theory of differential equations [9, 10, 11]. This re-emerging and fascinating topic is considered, by T. Padmanabhan, one of the Sleeping Beauties of theoretical physics [12].
Since in quantum mechanics frequency is related to energy, we could expect that the classical isoperiodicity shared by classical sheared potentials would translate into the property of quantum sheared potentials displaying the same energy spectra [15, 14, 13]. This is correct at the semi-classical level, but not in the full quantum theory [17, 16] (except in very few particular cases). In contrast with semi-classical theory, under shearing the spacing between consecutive energy levels can increase or decrease with the shearing parameter, and the eigenfunctions of Schrödinger’s equation,
$\left[\dfrac{\hbar^{2}}{2m}\dfrac{\partial^{2}}{\partial x^{2}}-\mathcal{U}_{\xi}(x)+\mathcal{E}\right]\psi_{\xi}(x)=0\,,$ (3)
can be substantially altered.
The process of shearing can be done by a single formula for all monomial wells, i.e., potentials of the form $\mathcal{U}(x)=k|x|^{n}$ by the relation
${\mathcal{U}}_{\nu}(x)=\begin{cases}\kappa(\nu|x|)^{n}&\text{when }x\geq 0\,,\\ \\ \kappa(\nu^{\prime}|x|)^{n}&\text{when }x<0\end{cases}$ (4)
where $\kappa$ is a positive constant and $\nu$ and $\nu^{\prime}$ are positive constants satisfying the shearing condition,
$\dfrac{1}{\nu}+\dfrac{1}{\nu^{\prime}}=2\quad\Longrightarrow\quad\nu^{\prime}=\dfrac{\nu}{2\nu-1}\,.$ (5)
We take $\nu\in(1/2,1]$ and accordingly $\nu^{\prime}\in[1,\infty)$. $\nu\rightarrow 1/2$ is the limit of an impenetrable barrier at the origin and $\nu=1$ is the case of symmetric potential. This is not the only way of shearing these families of potentials. In fact, there are infinite ways of shearing a particular potential.
In the next sections, we develop a method to study sheared potentials in quantum theory. We find the spectra and the exact eigenfunctions of the first two monomial wells (that is, the linear one and the harmonic oscillator). Usual analysis on sheared potentials focus only in the spectrum of energy. Indeed, the spectra show us that in the full quantum regime isospectrallity among potentials in the same sheared family is broken. In this paper we argue that studying the eigenfunctions allows us to see why this happens. By varying the shearing parameter the eigenfunctions changes continuously. This dance of the sheared eigenfunctions sheds light on what is happening to the spectrum.
2 Sheared Monomial Well of Order 1
Taking the case $n=1$ on formula (4) we get
${\mathcal{U}}_{\nu}(x)=\begin{cases}\kappa\nu x&\text{when }x\geq 0\,,\\ \\ -\kappa\left(\dfrac{\nu}{2\nu-1}\right)x&\text{when }x<0\end{cases}$ (6)
In Fig. 2 we plot three sheared potentials of this family.
$\mathcal{X}$$\mathcal{U}_{1/2}(x)$$\mathcal{U}_{1}(x)=k|x|$$\mathcal{U}_{4/7}(x)$
Figure 2: Three sheared potentials of the family of the first monomial well. The shearing condition is also indicated in the figure with the horizontal double arrows.
To solve Schrödinger’s equation with this potential, we suppose a piecewise solution
$\psi_{\nu}(x)=\begin{cases}\psi_{\nu}^{(+)}(x)\quad\text{when}\quad x\geq 0,\\ \psi_{\nu}^{(-)}(x)\quad\text{when}\quad x<0.\end{cases}$ (7)
which must satisfy the boundary conditions
$\psi_{\nu}^{(+)}(0)=\psi_{\nu}^{(-)}(0)\quad\text{and}\quad\dfrac{d\psi_{\nu}^{(+)}}{dx}(0)=\dfrac{d\psi_{\nu}^{(-)}}{dx}(0)\,.$ (8)
For $\nu=1$, we have the symmetric case, and the general solution is a combination of Airy functions [18],
$\psi_{1}{(x)}=\alpha\mathrm{Ai}\left(-\dfrac{E-kx}{k^{2/3}}\right)+\beta\mathrm{Bi}\left(-\dfrac{E-kx}{k^{2/3}}\right)\,,$ (9)
where we have defined $E=2m\mathcal{E}/\hbar^{2}$ and $k=2m\mathcal{\kappa}/\hbar^{2}$.
It is not difficult to generalize the previous result for an arbitrary $\nu$. In this case, it is straightforward to show that $\psi^{(\pm)}_{\nu}(x)$ are also given by sums of Airy functions,
$\psi_{\nu}^{(+)}(x)$ $=\alpha_{+}\mathrm{Ai}\left(-\dfrac{E-k\nu x}{(k\nu)^{2/3}}\right)+\beta_{+}\mathrm{Bi}\left(-\dfrac{E-k\nu x}{(k\nu)^{2/3}}\right)\,,$ (10)
$\psi_{\nu}^{(-)}(x)$ $=\alpha_{-}\mathrm{Ai}\left(-\dfrac{E+k\nu x/(2\nu-1)}{(k\nu/(2\nu-1))^{2/3}}\right)+\beta_{-}\mathrm{Bi}\left(-\dfrac{E+k\nu x/(2\nu-1)}{(k\nu/(2\nu-1))^{2/3}}\right)\,.$ (11)
The terms containing $\mathrm{Bi}$ must be discarded, since they diverge for $|x|\to\infty$. Using the conditions written in Eqs. 8, we find the following system of equations for the variables $\alpha_{+}$ and $\alpha_{-}$,
$\alpha_{+}\mathrm{Ai}\left(-\dfrac{E}{(k\nu)^{2/3}}\right)=\alpha_{-}\mathrm{Ai}\left(-\dfrac{E}{(k\nu/(2\nu-1))^{2/3}}\right)\,,$ (12)
$\alpha_{+}\left(k\nu\right)^{1/3}\mathrm{Ai}^{\prime}\left(-\dfrac{E}{(k\nu)^{2/3}}\right)=\alpha_{-}\left(\dfrac{k\nu}{2\nu-1}\right)^{1/3}\mathrm{Ai}^{\prime}\left(-\dfrac{E}{(k\nu/(2\nu-1))^{2/3}}\right)\,.$ (13)
From Eq.(12), we have
$\alpha_{-}=\alpha_{+}{\mathrm{Ai}\left(-\dfrac{E}{(k\nu)^{2/3}}\right)}{\mathrm{Ai}\left(-\dfrac{E}{(k\nu/(2\nu-1))^{2/3}}\right)}^{-1}\,.$ (14)
Using the previous equation, we obtain the explicit expressions for the wave functions,
$\psi_{\nu}^{(+)}(x)$ $=$ $\alpha_{+}\mathrm{Ai}\left(-\dfrac{E-k\nu x}{(k\nu)^{2/3}}\right)$ (15)
$\psi_{\nu}^{(-)}(x)$ $=$ $\alpha_{+}{\mathrm{Ai}\left(\!\!-\dfrac{E}{(k\nu)^{2/3}}\right)}{\mathrm{Ai}\!\!\left(\!\!-\dfrac{E}{(k\nu/(2\nu-1))^{2/3}}\right)}^{-1}\!\!\!\mathrm{Ai}\!\left(\!\!-\dfrac{E+k\nu x/(2\nu-1)}{(k\nu/(2\nu-1))^{2/3}}\right),$ (17)
where $\alpha_{+}$ is a mere normalization constant.
As it happens, the system formed by Eqs. (12) and (13) is homogeneous, meaning that it has non trivial solutions for $\alpha_{\pm}$ (i.e., solutions allowing $\alpha_{\pm}\neq 0$) if and only if the determinant of the coefficients vanish, leading to a transcendental equation for $E$
$\begin{split}&\left(-\dfrac{1}{2\nu-1}\right)^{1/3}\mathrm{Ai}\left(-\dfrac{E}{(k\nu)^{2/3}}\right)\mathrm{Ai}^{\prime}\left(-\dfrac{E}{(k\nu/(2\nu-1))^{2/3}}\right)\\ &=\mathrm{Ai}^{\prime}\left(-\dfrac{E}{(k\nu)^{2/3}}\right)\mathrm{Ai}\left(-\dfrac{E}{(k\nu/(2\nu-1))^{2/3}}\right)\,.\end{split}$ (20)
For a given value of $\nu$, which means a given potential well of the sheared family under consideration, the above equation is satisfied only for specific values of $E$, which are the allowed eigenvalues of the problem. In other words, finding the spectrum of the potential $\mathcal{U}_{\nu}(x)$ means determining the zeros of the function
$\begin{split}F_{\nu}(E):&=\left(-\dfrac{1}{2\nu-1}\right)^{1/3}\mathrm{Ai}\left(-\dfrac{E}{(k\nu)^{2/3}}\right)\mathrm{Ai}^{\prime}\left(-\dfrac{E}{(k\nu/(2\nu-1))^{2/3}}\right)\\ &-\mathrm{Ai}^{\prime}\left(-\dfrac{E}{(k\nu)^{2/3}}\right)\mathrm{Ai}\left(-\dfrac{E}{(k\nu/(2\nu-1))^{2/3}}\right)\,.\end{split}$ (21)
We solve this equation numerically.
Once we know how to calculate the energy spectrum of any potential well belonging to this family, we can investigate what happens to the energy levels as we vary continuously the parameter $\nu$. Let us denote by $E_{n}(\nu)$ the energy level $n$ of the potential $\mathcal{U}_{\nu}(x)$. In Fig. 3 we plot the first three energy levels as functions of $\nu$. In order to better visualize the changes in these energy levels during the shearing process, we plot each $E_{n}(\nu)$ normalized by $E_{n}(\nu=1)$ (the corresponding energy value for the symmetric potential of the family). Note that with this normalization, $E_{n}(\nu)/E_{n}(1)\rightarrow 1$ as $\nu\rightarrow 1$. From Fig. 3 we also see that, the larger the $n$, the smaller the impact on the energy spectrum of the shearing process. This was expected, since for large $n$ semi-classical results should hold and, as we have already mentioned, in the semi-classical limit the energy spectrum is $\nu$-independent. For this family, it seems that $n=4$ is already very large.
Figure 3: In the left panel we plot the first three normalized energy levels, $E_{n}(\nu)/E_{n}(1)$, with $n=0,1,2$, in terms of $\nu$. In the right panel we give a zoom in the plot of $E_{3}(\nu)$ in terms of $\nu$ to emphasize the oscillations for $\nu$ close to $1$.
As the title of this paper suggests, it is also very interesting to investigate the shape of the sheared eigenfunctions. The wavefunctions get progressively “expelled” from the negative half-line as $\nu\to 1/2$, until being totally restricted to the positive half-line as the impenetrable barrier limit is achieved. It is important to notice that the energies of each of these functions are different, as $E$ is $\nu-$dependent.
Figure 4: Complete eigenfunctions $\psi_{\nu}(x)$ for some values of $\nu$ for the first two energy levels (the fundamental state on the left and the first excited on the right). In this plots, we considered $k=1$.
We may connect some properties presented by the eigenfunctions with the spectra presented before. If we think of shearing as a physical process by which a given potential gets deformed, then it surely has an energy cost. An external agent is needed to provide the energy for shearing the potential under consideration and this is clearly seen in the $\nu$-dependence of the energy levels. Naturally, if the shearing is conducted abruptly it is well known that the quantum particle can make a transition to a different energy level. In order to analyse the behaviour of the eigenfunction we can think of an experiment where the system is prepared in a given eigenstate and the potential is very gently sheared, so that the particle remains on that state in the course of the whole process. In that picture, the particle would be “dragged” along with the deformation of the potential, meaning that the wavefunction describing the quantum particle is displaced through space and that the external force is performing work. At this point, it would be wise to bypass the disclaimer that one should exercise a good deal of caution when discussing forces in a quantum mechanical context. For our purposes, we can think of a force as an operator defined, in space representation, by $F(x)=-\mathcal{U}^{\prime}(x)$. As can be seen in eqs. (6), the force is proportional to the identity operator in this case, and given by $F_{+}=-\kappa\nu$ in the region $x\geq 0$ and by $F_{-}=\kappa\nu/(2\nu-1)$ in the region $x<0$, with the signs compatible with the restoring character of the force.
We show now that force can be a useful concept in order to understand the observed decrease of the energy eigenvalues near $\nu=0.5$ in Fig. (3). The key aspect is to observe that close to $\nu=0.5$ we have $|F_{-}|\gg|F_{+}|$. As we increase $\nu$, the wavefunction enters into the region $x<0$ being shifted to the left (and also deformed). In the process, a negative work is exerted in the region $x<0$ (where the force opposes the shift) while the opposite happens in the region $x>0$. Although the amount of probability that enters region $x<0$ is the same as that which leaves the region of positive $x$, the negative work overcomes the positive one due to the fact that $|F_{-}|\gg|F_{+}|$. This explains why $d\mathcal{E}/d\nu<0$ for $\nu$ close to $0.5$. When increasing $\nu$, the forces tend to equalize as we are approaching the symmetric case, and thus the work performed becomes very sensible to the deformation of the wavefunction and to its displacement for every $x$. This explains why the variations of energy are much smaller when $\nu$ is close to $1$ and means that we cannot anticipate the sign of the work and therefore of the energy change.
3 Sheared Monomial Well of Order 2 (Harmonic Oscillator)
The shearing of the harmonic oscillator is important due to its deep connection with isochronous potentials [15, 4, 5, 10], and for the central role that harmonic oscillators have in physics. This problem was already addressed in the context of problems associated to the semi-classical quantization [19], with a similar shearing rule we use in this paper. Here we investigate how the energy varies with the parameter $\nu$ so that we can quantify the breaking of isochronicity, showing the dance of the corresponding eigenfunctions in the process.
Taking the same steps of the previous case, we solve Schrödinger’s equation for the potential family
$\mathcal{U}_{\nu}(x)=\begin{cases}\kappa\nu^{2}x^{2}&\text{when }x\geq 0\,,\\ \\ \kappa{\dfrac{\nu^{2}}{(2\nu-1)^{2}}}x^{2}&\text{when }x<0\,.\end{cases}$ (22)
As we are dealing with the harmonic oscillator, is convenient to define $\kappa=m\omega^{2}/2$. The general solution of Schrödinger’s equation for the harmonic oscillator is (for $\nu=1$)
$\psi(x)=\alpha D_{\frac{\mathcal{E}}{\hbar\omega}-\frac{1}{2}}\left(\left(\dfrac{2m\omega}{\hbar}\right)^{1/2}|x|\right)+\beta D_{-\frac{\mathcal{E}}{\hbar\omega}-\frac{1}{2}}\left(i\left(\dfrac{2m\omega}{\hbar}\right)^{1/2}|x|\right)\,,$ (23)
where $D_{\alpha}(y)$ is the $D-$parabolic cylinder function [20]. When $y$ is purely imaginary, the function diverges for $|y|\to\infty$ so we can henceforth discard the second term in Eq. (23). For the sheared case, the wave functions are
$\psi_{\nu}^{(+)}(x)$ $=\alpha_{+}D_{\sigma(\nu)}\left(k\nu|x|\right)\,,$ (24)
$\psi_{\nu}^{(-)}(x)$ $=\alpha_{+}\left[\dfrac{D_{\sigma(\nu)}(0)}{D_{\sigma(\mu)}(0)}\right]D_{\sigma(\mu)}\left(k\mu|x|\right)\,,$ (25)
where $E=\dfrac{\mathcal{E}}{\hbar\omega}$, $k=\sqrt{2m\omega/\hbar}$, $\sigma(\nu)=\dfrac{E}{\nu}-\dfrac{1}{2}$, $\mu=\dfrac{\nu}{2\nu-1}$ and $\alpha_{+}$ is a mere normalization constant. Taking the same steps as the development for the linear potential, the transcendental equation for the spectrum in this case reads
$(2\nu-1)D_{\sigma(\nu)}^{\prime}(0)D_{\sigma(\mu)}(0)+D_{\sigma(\nu)}(0)D_{\sigma(\mu)}^{\prime}(0)=0\,.$ (26)
When $\sigma(\nu)$ is a positive integer, the $D-$parabolic cylinder function becomes proportional to a Hermite polynomial, as expected for the traditional harmonic oscillator.
Using expression (26), we can find the energy levels as function of the shearing parameter. The dependence on energy is, in fact, on $\sigma(\nu)$, and we recover the traditional spectrum of the harmonic oscillator with $\sigma(1)=n\in\mathds{N}$, ergo $E=(n+1/2)$. As we vary $\nu$, the first three energy levels change as shown in Figure (5). Here, we again have normalized our plots by $E_{n}(1)$ (the energy for the symmetric case).
Figure 5: In the left panel we plot the first three normalized energy levels for the sheared family of the harmonic oscillator, $E_{n}(\nu)/E_{n}(1)$, with $n=1,2,3$, in terms of $\nu$. In the right panel we give a zoom in the plot of $E_{2}(\nu)$ to emphasize the oscillations for $\nu$ close to $1$.
As we have seen in the first monomial well case, the impact of shearing is progressively diminished as $n$ gets larger, which is expected as we are approaching the semi-classical regime.
Now we analyse the behaviour of the eigenfunctions as we shear the potential. Naturally, they must vanish in the region $x<0$ as $\nu\to 1/2$.
Figure 6: Complete eigenfunctions $\psi_{\nu}(x)$ for some values of $\nu$ for the first two energy levels (the fundamental state on the left and the first excited on the right). In this plots, we considered $k=1$.
It is worth noticing that the eigenfunctions maintain their number of nodes for each energy levels, as it can be seen on Figures (4) and (6). This is, in fact, a general feature of one-dimensional Schrödinger’s equation [21], and to maintain the number of nodes, the eigenfunctions have to “drift” as they are sheared. Because of this, it was already expected that the form of the sheared eigenfunctions must have similar shape of the ones in the first monomial well, and this is expected for all of them.
We can again invoke the force concept in order to interpret some features connecting the dance of the eigenfunctions during the shearing process with the changes in the spectrum. The arguments are entirely similar and here we only remark that the relative change in energy is smaller than for the linear potential. This is expected since now the force vanishes at origin, reducing the work while the wavefunctions enter the region $x<0$.
This method can be used to treat other interesting potentials, the tricky part being the ability to solve the Schrödinger’s equation in both positive and negative semi-axis, i.e., to determine $\psi^{(\pm)}(x)$. Once the eigenfunctions are discovered, the sheared spectra and the sheared eigenfunctions can be obtained in the exact same fashion.
4 Conclusions and final remarks
In this work, we considered two families of sheared potential wells: a family of linear monomials and a family of harmonic oscillators. In both cases, we have chosen sheared potentials that retained the same functional form as the initial potential for $x<0$ and $x\geq 0$, but with different coefficients. For both families, we computed not only the spectra but also the corresponding eigenfunctions. In order to compute these spectra it was necessary to solve transcendental equations with the aid of numerical methods. According to Felix Klein (1849-1925), “it is well known that the central problem of the whole of modern mathematics is the study of transcendental functions defined by differential equations[22]. Although this modern is something old, the pursuit of this kind of treatment, combined with the tools of special functions and modern computing software, is a reviving topic in physics and mathematics on the last few decades [23].
As expected, we verified that the energy spectrum did not remain constant during the shearing process. We showed that this is due to the variation of the eigenfunctions. The main message of our work is that analysing the sheared eigenfunctions is a key element to understand why the spectra change. Combining the eigenfunctions with considerations about the work performed we could understand why the spectra display a decrease in energy with increasing $\nu$ in the vicinity of $\nu=0.5$, and also why the variations in this region is more significant then around $\nu=1$. This helps us understand the close similarities displayed by the results for the linear and the harmonic wells and we believe that these features would also be present for other power-law potentials. A perspective of this paper is to deepen this connection, evaluating quantitatively the work done in the process.
Once this calculation is made, one could investigate what would be the energetic costs during other processes of changing the potential wells in order to find the one which would require a minimum work to link the initial and final potentials. This kind of investigation could be useful to establish a connection with the promising and effervescent area of shortcut to adiabaticity.
Acknowledgments
All authors thank M. Asorey for insightful discussions. We also thank the Brazilian agencies CAPES, CNPq and FAPERJ for financial support. J.O.-C. is grateful to CAPES (Grant No. 88887.958122/2024-00). C.F. acknowledges funding from CNPq (Grants No. 308641/2022-1 and 408735/2023-6) and FAPERJ (Grant No. 204.376/2024). F.S.S.R. is grateful to CNPq (Grant No. 315730/2021-8).